# A generalized statistical model for fits to parton distributions

### v1

@@ IV.2 The GMM method compared to different tolerance criteria
Estimations of PDF uncertainties are needed for inference from QCD experiments at the LHC and other facilities. Various factors that determine the accuracy of PDFs include the theoretical uncertainty, the insufficiently understood source of uncertainty due to the prescription for experimental systematic errors, the tension between experimental data, and the sampling of PDF functional forms and fitting methodologies, as detailed in Ref. [35]. Global PDF analysis groups are aware of the shortcomings of using a fixed tolerance $\Delta\chi^{2}=1$ and have proposed alternate methods. Specifically, CTEQ-TEA [1] and MSHT [2] global PDF analyses, both adopt certain tolerance criteria, known as the two-tier tolerance and the dynamic tolerance criteria, respectively, on top of the basic Hessian method.
In Sec. IV.1, we compared results of the GMM with the usual LS method of determining 1-$\sigma$ contours by setting $\Delta\chi^{2}=1$. In this subsection, we compare the GMM results presented in Sec. IV.1 to the Hessian results with the CT and MSHT tolerance criteria.
We remind the reader that CTEQ-TEA Hessian eigenvector PDFs are constructed using a two-tier prescription [36, 37] that prevents unexpectedly large increments to either the global $\chi^{2}$ or to the $\chi^{2}_{E}$ values of individual experiments. Hence, the CT PDF error sets are determined using a two-tier (global+dynamic) tolerance ($T$) that accounts for theory, experimental, parametrization, and methodological uncertainties and results in error bands that are wider than the $\Delta\chi^{2}=1$ criterion. The CT Hessian uncertainty, at the 68% confidence level, nominally corresponds to $T^{2}\equiv\Delta\chi^{2}_{total}\approx 37$, where $\chi^{2}_{total}$ is the sum of the global $\chi^{2}$ and the so-called Tier-2 (dynamic) penalty  [36, 37]. The MSHT PDF fits by default apply the “dynamic tolerance” procedure, as detailed in [38, 39, 2]. This enlarges the uncertainties beyond the $\Delta\chi^{2}=1$ definition to account for data sets with tension, as well as potential mismatch of data and theory due to imprecision in the theory or parametrization and experimental measurements. The impact of different tolerance criteria on the determination of PDF induced uncertainties has been examined in Ref. [40]. In this study, we compare the results of applying the CT and MSHT tolerance criteria with $T^{2}=37$ to the ones with $\Delta\chi^{2}=1$.
@@
  • Case-1: Both Hessian sets with CT and MSHT tolerance criteria estimate a smaller pseudo-PDF uncertainty for $0.110^{-3}<x<$ 0.1, compared to the GMM. Recall for Case-1, the GMM produced a more faithful representation of uncertainties and overlapped with SG-A and SG-B. We see that the alternate tolerance criteria is not able to account for the larger uncertainties in this $x$ range. Further, the uncertainties given by the modified tolerance criteria are much larger than required when $x>0.1$. The same can also be observed in the fit parameter space. Here, the SG-C correlation ellipses with tolerance cover larger regions than the default Hessian set, but still cannot predict a correlation that spans regions of parameter space that cover portions of SG-A and SG-B.
@@ V Consistency checks for the GMM
In Sec. IV, we demonstrated how the GMM with two base Gaussian distributions is able to successfully describe the combined likelihood of two pseudo-data sets with tension. In this section, we show how the GMM reduces to the usual single Gaussian LS method when data are consistent. Further, as with any present day machine learning technique, when implementing the GMM, one must deal with over-fitting. In this section, we also remark on the consequences of over-fitting with the GMM.
@@ V.1 Consistent results from consistent data
  • Case-4: This case concerns a fit to the pseudo-data #$\#~{}7$, which has the same truth as pseudo-data #$\#~{}1$, but without any statistical fluctuations. In this case, only a single extremum of the log-likelihood of the GMM exists, i.e. $\theta_{i}=\theta_{k}$, and we shall see that numerically as well, the GMM will reduce to the LS method fit.
  • Case-5: This case concerns a fit to the pseudo-data #$\#~{}1$. Here, due to overfitting, we will see that although $\theta_{i}\simeq\theta_{k}$, the GMM tends to find multiple possible values of $\theta_{k}$. This in turn implies that the results from the GMM method are different from the LS method.
@@
The GMM method looks for tension and tries to categorize data into two or more different Gaussians. In general, for the GMM, overfitting gives rise to larger uncertainties and in that sense the GMM always provides a conservative estimate of the uncertainty. [^8] This is in contrast to overfitting by increasing the number of fit parameters in a LS fit, where overfitting may lead to a reduction in $\chi^{2}/N_{\text{pt}}<1$. In this sense, the GMM method provides conservative estimates of the uncertainty. However, one would like to provide a precise PDF which is essential for the physics goals of the community. In order to determine if there is overfitting in the GMM, we describe certain information criteria in the next section that can be used to determine the optimal number of Gaussians, i.e., the value of $K$. The result of Case-5 suggest that applying the GMM to a consistent data set will lead to larger uncertainties, which although conservative, should also be avoided.
@@
V.2 Information Criteria for an Optimal number of base Gaussians
The results shown in Sec. V.1 indicate that the GMM method might over-estimate the uncertainty, especially if one includes more base Gaussian distributions than needed. Therefore, a procedure to determine the optimal value of $K$, for a mixture of $K$ base Gaussian distributions of the GMM, is required. Here, we propose the use of the Akaike information criterion (AIC) [41] and Bayesian information criterion (BIC) [42] as indicators that will help determine optimal values for $K$, i.e., the number of Gaussians. The values of AIC and BIC, defined in Eq. (25), increase as more parameters are introduced. They also take into account the change in goodness of fit and reduce as the fit improves. Both AIC and BIC are thus able to account for the trade-off between introducing additional parameters and an improvement to the fit. The minimum values for the AIC and BIC indicate to us the best choice of parameters, in this case, the best choice for the value of $K$. The AIC and BIC are defined as
@@ V.2 Information Criteria for an Optimal number of base Gaussians
$N_{p}=K\times N_{\text{parm}}+(K-1),$ (26)
@@
$N_{\text{parm}}=2K+(K-1).$ (27)
We scan the AIC and BIC of the GMM up to $K=4$ for all 5 cases considered in this work, as summarized in Table 5. For Case-1 and Case-3, both AIC and BIC suggest two base Gaussian distributions would be optimal in order to faithfully extract information from their pseudo-data sets. This confirms the results shown in Sec. IV, particularly in Figs. 4 and 6, where two base Gaussian distributions can estimate a correlation that covers the two pseudo-data sets in tension. The AIC and BIC suggest for Case-2 a sinlge Gaussian distribution is enough to fit this pseudo-data set. In Fig. 5, the LS method predicts a correlation, whose slope is close to the correlation estimated with two base Gaussian distributions. This is consistent with the fact that Case-2 has minimal tension due to the large values of data errors. At this point we would like to point out that the GMM in conjunction with AIC and BIC, can be used to determine how much the errors in the data should be inflated in order to make them consistent so that we can use the LS method, as suggested in Ref. [16, 27, 26, 28, 29].999 Specifically, one possible measure is to increase the errors in the data sets until the AIC and BIC indicate that $K=1$ is the optimal value of $K$. This provides an alternate way to determine the scale factor used in Ref. [26]. For Case-4 and Case-5, the AIC and BIC of course suggest that there is no need of using the GMM method, since by doing so one is over-fitting the data. We reiterate that the consequence of over-fitting is to increase our uncertainty and therefore leads to a conservative estimate of uncertainty, i.e., using the optimum value of $K$ provides the most precise PDFs, whereas larger values of $K$ will lead to PDFs with larger uncertainty. We also notice in Table 5 that a higher number of base distributions $K$ sometimes would lead to equal or slightly lower value of $-\ln L$ than the $K$ suggested by AIC and BIC, indicating that one can equally or even better fit the data with more base distributions. However this is likely due to over-fitting the statistical fluctuations in the data. To conclude, although the GMM method is expected to reduce to the LS method for a consistent data set, one should avoid using it without detailed assessment of the level of tension. The combination of AIC and BIC indicators with the generalized statistical model of the GMM method could effectively reproduce the correct correlation in the presence of inconsistency in data. At the same time we can determine optimal values of $K$ and provide a formalism to determine more precise PDFs.
@@
$N_{\text{pt}}$=50 $-\ln L$88.6 -88.6 -88.6 -88.6
Table 5: The AIC, BIC and $-\ln L$ for Case-1 to Case-5 obtained by the GMM with various numbers of base distributions. The lowest values of AIC, BIC and $-\ln L$ for each cases are written in bold font.
@@ VI Conclusion
An accurate PDF uncertainty estimation is a theoretical necessity for precision measurements as well as new physics searches at the LHC and future colliders. The sources of PDF uncertainty include the theoretical uncertainty, the experimental systematic errors, and the sampling of PDF functional forms and fitting methodologies, etc. [35]. The existence of tension among different data sets can also reduce the precision of PDFs and necessitates the use of alternate statistical models that provide more accurate estimates of uncertainty. In this work, we described a formalism that generalizes the commonly used Least-Squares (LS) method and provides a way to combine likelihoods from different experiments in a statistically robust and meaningful way.
The proposed generalized likelihood modeled via the Gaussian mixture model (GMM) includes the possibility of the likelihood being either multi-modal or uni-modal. We demonstrated the shortcomings of the LS model when dealing with data sets that are in tension, using a toy model of PDFs. We found that, with a single Gaussian distribution assumed in the LS method, neither the Monte Carlo method nor the Hessian method can provide an accurate sampling of the probability distribution of the pseudo PDFs. Although the Monte Carlo method directly samples the probability distribution in the data space, the assumption that there is only one Gaussian distribution in the PDF space leads to a LS fit that would be stuck in a narrow valley of $\chi^{2}$. The same situation applies to the Hessian method. Inflating the Hessian uncertainty with certain tolerance criteria is a conservative prescription for the uncertainty estimation currently used by PDF fitting collaborations. We find that doing so is not a guarantee of a good approximation of the combined likelihood of two data sets that are in tension, particularly in the case where the tension is complicated. Specifically, the value of the tolerance chosen is not based on statistically robust arguments. Further, we showed how increasing the tolerance does not accurately capture tension and leads to inadequate uncertainty in some regions of the PDF but at the same time an overestimation of uncertainty in other regions of the PDF, c.f. Fig. 4. This can also be seen in the correlation ellipses of the fit parameter space. What is required here is not just an extension of the original $\chi^{2}$ ellipse, but also a rotation of the axes of the correlation ellipse. The latter can neither be achieved with a modified tolerance criteria nor by using a Monte-Carlo sampling technique.
@@
The novel GMM pseudo-PDF fit overcomes these shortcomings of the LS method by parametrizing the likelihood as a mixture of Gaussian distributions, by allowing for a multi-modal feature of the likelihood. We have described how this unsupervised machine learning technique can be leveraged to learn about tension in the fits as well as be used to determine precisely the uncertainties on the PDFs. Specifically, it is possible to use the GMM for performing the unsupervised machine learning task of clustering the data in order to identify and learn about the tension in the data sets. We have demonstrated how the GMM is simply generalization of the $\chi^{2}$ log-likelihood and reduces to it when data is consistent. Finally, It is possible to over-fit in the GMM, where the model might end up learning statistical fluctuations rather than real tension between data sets. We have described the use of the Akaike information criterion (AIC) and Bayesian information criterion (BIC) to avoid over-fitting and can be used to determine optimal value of the number of base distributions for the GMM method.

### v2

@@ IV.2 The GMM method compared to different tolerance criteria
Various factors that determine the precision of PDFs include theoretical uncertainties, tension between experimental data, and the sampling of PDF functional forms and fitting methodologies, often referred to as epistemic uncertainties, c.f. Ref. [44].
Global PDF analysis groups are aware of the shortcomings of using a fixed tolerance $\Delta\chi^{2}=1$ and have proposed alternate methods. Specifically, CTEQ-TEA [1] and MSHT [2] global PDF analyses, adopt certain tolerance criteria, known as the two-tier tolerance and the dynamic tolerance criteria, respectively, on top of the basic Hessian method.
In Sec. IV.1, we compared results of the GMM with the usual LS method of determining 1-$\sigma$ contours by setting $\Delta\chi^{2}=1$. In this subsection, we compare the GMM results presented in Sec. IV.1 to the Hessian results with the CTEQ-TEA and MSHT tolerance criteria.
We remind the reader that CTEQ-TEA Hessian eigenvector PDFs are constructed using a two-tier prescription [45, 46] that prevents unexpectedly large increments to either the global $\chi^{2}$ or to the $\chi^{2}_{E}$ values of individual experiments. Hence, the CTEQ-TEA PDF error sets are determined using a two-tier (global+dynamic) tolerance ($T$) that accounts for theory, experimental, parametrization, and methodological uncertainties and results in error bands that are wider than the $\Delta\chi^{2}=1$ criterion. The CTEQ-TEA Hessian uncertainty, at the 68% confidence level, nominally corresponds to $T^{2}\equiv\Delta\chi^{2}_{total}\approx 37$, where $\chi^{2}_{total}$ is the sum of the global $\chi^{2}$ and the so-called Tier-2 (dynamic) penalty  [45, 46]. The MSHT PDF fits by default apply the “dynamic tolerance” procedure, as detailed in [47, 48, 2]. This enlarges the uncertainties beyond the $\Delta\chi^{2}=1$ definition to account for data sets with tension, as well as potential mismatch of data and theory due to imprecise theory or parametrization. The impact of different tolerance criteria on the determination of PDF induced uncertainties has been examined in Ref. [49]. In this study, we compare the results of applying the CTEQ-TEA and MSHT like tolerance criteria with $T^{2}=37$ to the ones with $\Delta\chi^{2}=1$.
@@
  • Case-1: Both Hessian sets with CTEQ and MSHT tolerance criteria estimate a smaller pseudo-PDF uncertainty for $10^{-3}<x<0.1$, compared to the GMM. Recall for Case-1, the GMM produced a more faithful representation of uncertainties and overlapped with SG-A and SG-B. We see that the alternate tolerance criteria is not able to account for the larger uncertainties in this $x$ range. Further, the uncertainties given by the modified tolerance criteria are much larger than required when $x>0.1$. The same can also be observed in the fit parameter space. Here, the SG-C correlation ellipses with tolerance cover larger regions than the default Hessian set, but still cannot predict a correlation that spans regions of parameter space that cover portions of SG-A and SG-B.
@@ V Consistency checks for the GMM
In Sec. IV, we demonstrated how the GMM with two base Gaussian distributions provides a better description of the combined likelihood of two pseudo-data sets with tension and provides more accurate uncertainties. In this section, we show how the GMM reduces to the usual single Gaussian LS method when data are consistent. Further, as with any present day machine learning technique, when implementing the GMM, one must contend with over-fitting. In this section, we also remark on the consequences of over-fitting with the GMM.
@@ V.1 Consistent results from consistent data
  • Case-4: This case concerns a fit to the pseudo-data #7\#~7,
    which is the same as pseudo-data #1\#~1, but without any statistical fluctuations.
    In this case, only a single extremum of the log-likelihood of the GMM exists, i.e. $\theta_{i}=\theta_{k}$, and we shall see that numerically as well, the GMM reduces to the LS method fit.
  • Case-5: Here, we fit only to a single data set, namely pseudo-data set #1\#~1. Due to overfitting, we will see that although $\theta_{i}\simeq\theta_{k}$, the GMM tends to find multiple possible values of $\theta_{k}$. This in turn implies that the results from the GMM method are different from the LS method.
@@
The GMM method looks for tension and tries to categorize data into two or more different Gaussians. In general, for the GMM, overfitting gives rise to larger uncertainties and in that sense the GMM always provides a conservative estimate of the uncertainty. However, one would like to provide a precise PDF which is essential for the physics goals of the community. In order to determine if there is overfitting in the GMM, we describe certain information criteria in the next section that can be used to determine the optimal number of Gaussians, i.e., the value of $K$. The result of Case-5 suggest that applying the GMM to a consistent data set will lead to larger uncertainties, which although conservative, should also be avoided.
@@
V.2 Determining the Number of Gaussian Modes Using Information Criteria
The results shown in Sec. V.1 indicate that the GMM method overestimates the uncertainty for consistent data sets, especially if one includes more base Gaussian distributions than needed.
Although the GMM automatically clusters data into sets with tension and those that are consistent, a method to quantify the level of consistency and to determine the optimal number of base Gaussian distributions ($K$) required to avoid overfitting and therefore overestimation of uncertainties.
Here, we propose the use of the Akaike information criterion (AIC) [50] and Bayesian information criterion (BIC) [51] as indicators that will help determine optimal values for $K$, i.e., the number of Gaussians, and as a test of the consistency of fits.
The values of AIC and BIC, defined in Eq. (23), increase as more parameters are introduced. They also take into account the change in goodness of fit and reduce as the fit improves. Both AIC and BIC are thus able to account for the trade-off between introducing additional parameters and an improvement to the fit. The smallest values of the AIC and BIC indicate to us the best choice of parameters, in this case, the best choice for the value of $K$. The AIC and BIC are defined as
@@ V.2 Determining the Number of Gaussian Modes Using Information Criteria
$N_{p}=N_{\text{parm}}^{\text{GMM}}=K\times N_{\text{parm}}+(K-1),$ (24)
@@
$N_{\text{parm}}^{\text{GMM}}\bigg{|}_{N_{\text{parm}}=2}=2K+(K-1).$ (25)
We scan the AIC and BIC of the GMM up to $K=4$ for all 5 cases considered in this work, as summarized in Table 3. For Case-1 and Case-3, both AIC and BIC suggest two base Gaussian distributions are optimal. This confirms the results shown in Sec. IV, particularly in Figs. 4 and 6, where two base Gaussian distributions can estimate a correlation that covers the two pseudo-data sets in tension. The AIC and BIC suggest a single Gaussian distribution suffices for Case-2. In Fig. 5, the LS method predicts a correlation, whose slope is close to the correlation estimated with two base Gaussian distributions. This is consistent with the fact that Case-2 has minimal tension due to the large values of data errors. At this point we would like to point out that the GMM in conjunction with AIC and BIC, can be used to determine how much the errors in the data should be inflated in order to make them consistent so that we can use the LS method, as suggested in Ref. [16, 24, 38, 52, 53].101010 Specifically, one possible measure is to increase the errors in the data sets until the AIC and BIC indicate that $K=1$ is the optimal value of $K$. This provides an alternate way to determine the scale factor used in Ref. [38], c.f. Appendix C.
For Case-4 and Case-5, the AIC and BIC, as expected, suggest that there is no need of using $K>1$, since by doing so one is over-fitting the data. We reiterate that the consequence of over-fitting is to increase our uncertainty and therefore leads to a conservative estimate of uncertainty, i.e., using the optimum value of $K$ provides the most precise PDFs, whereas larger values of $K$ will lead to PDFs with larger uncertainty. We also notice in Table 3 that a higher number of base distributions $K$ sometimes leads to equal or slightly lower values of $\ln L$.
Thus, if we were to only use $\ln L$ as the measure of the goodness of fit, and not use the AIC and BIC, we would end up overfitting and choosing larger values of $K$.
To conclude, although the GMM method is expected to reduce to the LS method for a consistent data set, one should avoid using it without detailed assessment of the level of tension. The combination of AIC and BIC indicators with the generalized statistical model of the GMM method could effectively reproduce the correct correlation in the presence of inconsistency in data. At the same time we can determine optimal values of $K$ and provide a formalism to determine more precise PDFs.
@@
$N_{\text{pt}}$=50 $\ln L -$88.6 -88.6 -88.6 -88.6
Table 3: The AIC, BIC and $\ln L$ for Case-1 to Case-5 obtained by the GMM with various numbers of base distributions. The lowest values of AIC, BIC and $\ln L$ for each cases are written in bold font.
@@ VI Conclusion
Accurate PDF uncertainty estimation is a theoretical necessity for precision measurements as well as new physics searches at the LHC and future colliders. Several sources of PDF uncertainty exist, including theoretical uncertainty, experimental systematic errors, and sampling of PDF functional forms and fitting methodologies, see for example [44]. The existence of tension among different experimental data sets can reduce the precision of PDFs and necessitates the use of alternate statistical models that provide more precise estimates of uncertainty. In the presence of tension, expansion of uncertainty is necessary to make fits consistent, and methods to do so have been proposed and used by various PDF fitting groups. The nature and amount of expansion of uncertainty differs for different methodologies available in literature, many of which introduce ad hoc criteria.
In this work, we have introduced a novel method to estimate uncertainties for inconsistent data sets that avoids the use of ad hoc criteria, checks consistency and produces expanded uncertainties in a statistically meaningful way. Specifically, for inconsistent data, uncertainties need to be expanded so that the central values of the separate fits to each consistent data set lies close to or within the uncertainty bands of the theory parameters for the combined fit. This ensures a more precise representation of the range of possible theory values, given the inconsistent data, and avoids biasing the combined fit. The proposed generalized likelihood, modeled via the Gaussian mixture model (GMM), is able to appropriately map uncertainties by including the possibility that the likelihood can be either multi-modal or uni-modal.
In this work, we demonstrated the shortcomings of the LS model when combining data sets that are in tension, using a toy model of PDFs. For LS fitting, the Monte Carlo method and the Hessian method provide a narrow uncertainty band that is not representative of the tension in the data sets. Increasing the uncertainty, either by scaling the uncertainty on each measurement or equivalently, by introducing a static tolerance criteria, does address the problem but is not very precise. For example, using such criteria may increase uncertainty in regions where it is not necessary to do so whereas the uncertainty in regions of parameter space that do have tension is not suitably adjusted. Hence, the tension is not appropriately represented by the uncertainty bands of the LS fit.
@@
The novel GMM pseudo-PDF fit overcomes these shortcomings of the LS method by parametrizing the likelihood as a mixture of Gaussian distributions and by allowing for a multi-modal feature of the likelihood. We have described how this unsupervised machine learning technique can be leveraged to learn about tension in the fits as well as be used to determine precisely the uncertainties on the PDFs. We have shown how to use the Akaike information criterion (AIC) and Bayesian information criterion (BIC) to avoid overfitting, determine the number of Gaussian modes of the GMM, and check for consistency of the fits. Specifically, it is possible to use the GMM for performing the unsupervised machine learning task of clustering the data in order to identify and learn about the tension in the data sets. Moreover, we have demonstrated how the GMM reduces to the usual LS method in the absence of tension. Finally, we have also provided a more intuitive Bayesian interpretation of the GMM method and discussed its connection to Ensemble Averaging and Bayesian Model Averaging.